OpenStudy (anonymous):

Please help! :) What are the solutions for x in the proportion x plus 6 all over 2 times x equals x over 2.?

OpenStudy (anonymous):

\[\frac{ x+6 }{ 2x }= \frac{ x }{ 2 }\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

yes that's the question, @k.rajabhishek

OpenStudy (helder_edwin):